OpenStudy (montsechamorro):

Please help. Passes through (-1,4) perpendicular to 3x+2y=9

OpenStudy (itz_sid):

Okay so you should start off by putting that equation into slope-intercept form. It will be easier to work with that way.

OpenStudy (jiteshmeghwal9):

Question is incomplete

OpenStudy (shadowlegendx):

You can assume its asking for a line with those parameters

OpenStudy (shadowlegendx):

As @iTz_Sid said, we must put the equation into slope-intercept form. \[3x+2y=9 \] Subtract 3x from both sides \[2y = -3x + 9\] Divide the entire equation by 2 \[y = \frac{ -3}{ 2 }x + \frac{ 9 }{ 2 }\]

OpenStudy (montsechamorro):

Oh! Let me do the problem really quick and see if we get the same answer

OpenStudy (shadowlegendx):

We need the slope perpendicular to -3/2x, so we take the opposite reciprocal. Switch the numerator(-3) with the denominator(2). We get 2/-3. Then we change the negative sign to a positive sign. The slope 2/3 is perpendicular to -3/2

OpenStudy (shadowlegendx):

Ok, I won't write any further steps :P

OpenStudy (montsechamorro):

Okay so I got y= 2/3x+4/3

OpenStudy (shadowlegendx):

How did you get 4/3

OpenStudy (shadowlegendx):

Oh as your final answer?

OpenStudy (montsechamorro):

Yes

OpenStudy (jiteshmeghwal9):

If u have an equation of format ax+by+c=0. Then a line perpendicular to this line shall be ax-by+c' where c' is another constant

OpenStudy (jiteshmeghwal9):

3x+2y-9=0 Perpendicular line is 3x-2y+c'=0 Replace x &y by -1 & 4 respectively to get c'

OpenStudy (jiteshmeghwal9):

Thus form the final equation

OpenStudy (montsechamorro):

So am I wrong with y=2/3x+4/3

OpenStudy (shadowlegendx):

Yes

OpenStudy (montsechamorro):

Alright so I used y-4=2/3(x-4)

OpenStudy (shadowlegendx):

Why x-4

OpenStudy (montsechamorro):

I used the y-y1=M(x-x1) formula

OpenStudy (montsechamorro):

oh I see! I plugged in 4

OpenStudy (shadowlegendx):

yeah c;

OpenStudy (montsechamorro):

So y=2/3x+21/12

OpenStudy (shadowlegendx):

woah

OpenStudy (shadowlegendx):

no

OpenStudy (shadowlegendx):

What did you start with?

OpenStudy (montsechamorro):

Y-4= 2/3(x+1)

OpenStudy (shadowlegendx):

Then what did you do next

OpenStudy (montsechamorro):

distribute 2/3 Y-4= 2/3x+ 2/3

OpenStudy (shadowlegendx):

Ok, then what next

OpenStudy (montsechamorro):

ADD 4 to 2/3

OpenStudy (montsechamorro):

oh! I found my mistake. So my answer was y= 2/3x+ 14/3

OpenStudy (montsechamorro):

is that right

OpenStudy (shadowlegendx):

Yes, sorry ate dinner

OpenStudy (montsechamorro):

No worries :) thank you so much